# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 3.1  Upper bound
We begin with a concept that appears in several previous approaches [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover]. For any $u\in\partial P$, let $v\in[u,2n+1)$ be the maximal value such that there exists a guard $(g,[u,v])$. This defines a function $\textnormal{{next}}:\partial P\to\partial P$ mapping $u$ to $v$. Let $k$ denote the minimum number of guards required to cover $\partial P$. If we choose any point $x\in\partial P$ and apply next a total of $k+1$ times, the resulting point lies beyond $x+n$. Our goal is to compute a set $X\subset\partial P$ such that for at least one $x\in X$ there exists an optimal solution of $k$ guards containing a guard $(g,[x,v])$. If we recursively apply the next function $k$ times to all points in $X$, this yields an optimal solution using $k$ guards. Hence, if $X$ has linear size and if we can evaluate next in $O(\log n)$ time, we obtain an $O(kn\log n)$-time algorithm.
Classifying guards.
Section˜4 introduces the structural results to compute $X$ and the next function. We define three collections of highly structured guards such that every guard is dominated by one of these three types. Specifically, we partition the set of all guards into good and bad guards. Intuitively, a guard $(g,[u,v])$ is good if the angle $\sphericalangle(u,g,v)$ is at most $\pi$, and bad otherwise. We then define three collections of dominators: the good, the bad, and the ugly.
@@
We can now provide the high-level intuition for why $\textnormal{{next}}(x)$ can be evaluated in logarithmic time. If we precompute all good and bad dominators $(g,[u,v])$, then we can store their corresponding intervals $[u,v]$ in a segment tree $T$. For any $x\in\partial P$, we performing a stabbing query on $T$ and we return the maximal right endpoint $r$ among the intervals stabbed by $x$. Let $v=\textnormal{{next}}(x)$. If there exists a guard $(g,[u,v])$ dominated by a good or bad dominator, then $v=r$. We then compute an ugly dominator for $x$ on the fly and compare its result to $r$ to obtain the correct value of $\textnormal{{next}}(x)$.
@@
[…]
The reduced good dominators $\mathcal{D}$ are not only essential for computing next, but they also induce the set $X$. Specifically, we prove that there always exists an optimal solution Opt to the Contiguous Art Gallery problem such that at least one guard $(g,[u,v])\in\textsc{Opt}$ either belongs to $\mathcal{D}$ or has $u$ as a vertex of $P$. We define $X$ as the set of all vertices of $P$, together with all points $u\in\partial P$ such that there is a reduced good dominator $(g,[u,v])$.
Sliding windows.
We use a sliding window over indices $(i,j)$ to compute (a superset of) $\mathcal{D}$. The remainder of our $O(n\log n)$-time algorithms rely on a different but very specific sliding window:
Definition 1.
@@
In Section˜6 we compute linear-size sets $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ where $\mathcal{D}\subset\mathcal{D}^{\prime}$ and $\mathcal{B}\subset\mathcal{B}_{\sigma}$. We first compute $\mathcal{D}^{\prime}$ using a sliding-window algorithm. Note that we cannot use a conforming sliding sequence $\sigma$ for this task, as a reduced good dominator may be strictly dominated by a guard that is not a good dominator. Consequently, the defining pair $(i,j)$ for a guard $(g,[u,v])\in\mathcal{D}$ need not belong to $\sigma$.
@@
For all subsequent algorithms, we use a conforming sliding sequence $\sigma$, which we compute in an online fashion starting from $(i,j)=(1,1)$ and repeatedly incrementing either $i$ or $j$. The choice of which index to increment follows a simple rule: while there exists a guard $(g,[u,v])$ such that $i,j+1]\subset[u,v]$, we increment $j$ and add the new pair $(i,j)$ to $\sigma$; otherwise, we increment $i$ and add the new pair $(i,j)$ to $\sigma$. By the result of [Chan2019Time-Windowed], we can maintain $\mathcal{E}[i-1,j+2]$ as a balanced binary tree while incrementing $i$ and $j$. We show that, given $i$, $j$, and $\mathcal{E}[i-1,j+2]$, this decision can be made in $O(\log n)$ time, yielding an overall $O(n\log n)$ algorithm for constructing $\sigma$. From $\sigma$, we derive a linear-size set $\mathcal{B}_{\sigma}$ containing $\mathcal{B}$. We store the visibility of guards in $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ in an segment tree $T$. Thus, we may obtain for any point $x\in\partial P$ the guard $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$ that has $x\in[u,v]$ and maximizes $v$.
An $O(kn\log n)$ algorithm.
@@
In Section˜7, we consider the set $X$ and iterate over all $(i,j)\in\sigma$. For each $x\in X$, we perform a stabbing query on $T$ to find the maximum value $r$ such that there exists a guard $(g,[x,r])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$. The value $\textnormal{{next}}(x)$ differs from $r$ only if it is realised by an ugly dominator. The value $\lfloor r\rfloor$ serves as a lower bound for the index $j$ such that $\textnormal{{next}}(x)\in(j,j+1]$, and by definition $(i,j)\in\sigma$. These observations yield an algorithm that iterates over $\sigma$ and computes $\textnormal{{next}}(x)$ for all $x\in X$ in $O(n\log n)$ total time. Repeating this process $k$ times produces an optimal solution in $O(kn\log n)$ time. This is already a major improvement over the state-of-the-art [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover] and serves as a crucial subroutine for our $O(n\log n)$-time algorithm.
An $O(n\log n)$ algorithm.
@@
We compute a representation of next as a piecewise, possibly degenerate, Möbius transformation. This approach was first proposed in [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover], where the resulting function consisted of $O(n^{3})$ pieces and was computed in $\tilde{O}(n^{5})$ time. In contrast, we use our dominators to define a function with only $O(n)$ pieces, which can be computed in $O(n\log n)$ time. Specifically, in Section˜8, we partition the boundary $\partial P$ into intervals $I$ such that either:
@@
3.2  Lower bound
In Section˜10, we reduce from Set Disjointness, where we are given two sets $A,B\subset[0,n^{3}]$, each of size $n$, and must decide whether $A\cap B=\emptyset$. While in the wordRAM-model there is a deterministic $O(n\log\log n)$-time algorithm for solving this problem (via sorting [Han2002Deterministic]), in comparison-based models of computation such as the realRAM-model, it has a worst-case lower bound of $\Omega(n\log n)$, even if $A$ is already sorted [YaoLowerBound]. Given $A,B\subset[1,n^{3}]$ with $A$ sorted, we can in linear time construct a simple polygon $P$ such that $A\cap B=\emptyset$ if and only if $P$ can be guarded with $2|B|+2$ guards (see Figure˜2).
We begin with a triangle $T$ whose top edges, $t_{1}:$ $(0,0)$ to $(\frac{n^{3}}{2},\frac{n^{3}}{2})$ and $t_{2}:$ $(\frac{n^{3}}{2},\frac{n^{3}}{2})$ to $(n^{3},0)$, represent the interval $0,n^{3}]$. We construct the polygon in cyclical clockwise order, starting from $(0,0)$. For each of the increasing values $a_{i}\in A$, we make an indent in $t_{1}\cup t_{2}$ excluding the point corresponding to $a_{i}$ from $P$. Next, we partition a small segment on the bottom edge ($(n^{3},0)$ to $(0,0)$) into $|B|$ intervals $I_{j}$. For each $b_{j}\in B$, we introduce a pocket: two consecutive indentations of $I_{j}$ such that a guard can see $I_{j}$ if and only if it stands at the point of $t_{1}$ or $t_{2}$ corresponding to $b_{j}$. Hence, there exists a guard that can cover $I_{j}$ if and only if $b_{j}\notin A$. Finally, we place a blocker between consecutive intervals $I_{j}$ and $I_{j+1}$ such that each $I_{j}$ and each blocker requires a unique guard. The resulting polygon $P$ can be constructed in linear time and guarded with $2|B|+2$ contiguous guards if and only if $A\cap B=\emptyset$, establishing an $\Omega(n\log n)$ lower bound for the Contiguous Art Gallery problem in comparison-based models of computation such as the realRAM-model.
@@ 4  A combinatorial classification of guards
We define three types of guards such that every guard is dominated by one of these three types. To this end, we first partition the set of all guards into good and bad guards. In Figure˜1, $g_{2}$, $g_{3}$, and $g_{4}$ are good, while $g_{1}$ and $g_{5}$ are bad.
Definition 2.
@@
We will define three collections of guards called dominators: the good, the bad, and the ugly. We prove that all good guards are dominated by a good dominator, which will induce a set $X$ of $O(n)$ points along $\partial P$ such that there exists an optimal solution that includes a guard $(g,[x,v])$ for $x\in X$. We prove that all bad guards are dominated by either a bad or an ugly dominator. The flowchart in Figure 3 illustrates our approach. To this end, we first make some observations:
Lemma 2.
@@
Proof.
By maximality of $v$, $g\neq v$ and there exists an $\varepsilon^{*}>0$ such that $\forall\varepsilon\in(0,\varepsilon^{*}]$ the shortest path from $g$ to $v+\varepsilon$ visits a reflex vertex $x$. If $x=v$ then $v$ is a reflex vertex of $P$. Otherwise, $\overline{v\,g}$ contains $x$ in its interior. ∎
Lemma 3.
@@
Proof.
For a guard that is inclusion-wise maximal it cannot be that $g=u$ or $g=v$. That the shortest path from $u$ to $v$ is left-turning follows from the fact that $g$ can see $u$ and $v$ and $\sphericalangle(u,g,v)>\pi$. If $u$ (respectively $v$) is not a vertex of $P$, then Lemma˜2 implies that the edge $\overline{u\,g}$ (respectively $\overline{v\,g}$) contains a reflex vertex of $P$ in its interior, which must appear as an interior vertex on the shortest path from $u$ to $v$.
If only $v$ is maximal, then it again cannot be that $g=v$, but it may be that $g=u$. In this case the shortest path from $u$ to $v$ is the segment $\overline{u\,v}$, which is also left-turning. If $v$ is not a vertex of $P$, Lemma˜2 still implies there is at least one vertex on this shortest path. ∎
Figure 3: Flowchart illustrating how a guard is dominated. We prove the green block in Lemma˜5 and the red block in Lemma˜6.
4.1  The good dominators
We show that there exists a set $\mathcal{D}$ of $O(n)$ guards, called the reduced good dominators, such that every good guard is dominated by one of them. We begin by defining two categories of good dominators, each of size $O(n^{2})$.
Observation 4.
@@ 4.1  The good dominators
Proof.
For any fixed point $g\in P$, the set of points in $P$ visible from $g$ forms the visibility polygon of $g$. Its intersection with $\partial P$ induces a fixed set of pairwise disjoint maximal chains, and only one of these chains can contain $[i,j]$. ∎
Definition 3.
@@ 1. $\mathbf{A}$
    (segment-intersecting dominator): See Figure˜5. If $\mathcal{E}[i-1,j+1]$ intersects the interior of $\overline{i\,j}$, we define the guard $(g,[u_{\text{max}},v_{\text{max}}])$, where $g$ is the last point of $\overline{i\,j}$ intersecting $\mathcal{E}[i-1,j+1]$, and $u_{\text{max}},v_{\text{max}}]\subset[i-1,j+1]$ is the maximal chain containing $[i,j]$ that is visible to $g$.
@@ 2. $\mathbf{B}$
    (angle-maximizing dominator): See Figure˜5. If $\mathcal{E}[i-1,j+1]$ lies left of $\overline{i\,j}$, we define the guard $(g,[u_{\text{max}},v_{\text{max}}])$. We define $g$ as $i$ (or $j$) if $\mathcal{E}[i-1,j+1]$ intersects $i$ (or $j$). Otherwise, $g$ is the point in $\mathcal{E}[i-1,j+1]$ that maximizes the angle $\sphericalangle(i,g,j)$. We define $u_{\text{max}},v_{\text{max}}]\subset[i-1,j+1]$ as its maximal visible chain containing $[i,j]$.
@@
The guard $(g,[u,v])$ is good and thus $g\neq u,v$ and $\sphericalangle(v,g,u)\leq\pi$ by Definition˜3. We prove that a good dominator (and thus a reduced good dominator) dominates $(g,[u,v])$. Let $u\in[i-1,i)$ and $v\in(j,j+1]$. If $i=j$ or i1=ji-1=j, then the dominator for $(i,j)$, which is the guard (i,[i1,j+1])(i,[i-1,j+1]), dominates $(g,[u,v])$. So, assume $i<j$.
@@
We define by $P_{g}$ the polygon formed by the edges $\overline{v\,g}$, $\overline{g\,u}$ and $[u,v]$. Since $(\overline{u\,g},\overline{v\,g})$ forms a convex wedge, and $P_{g}$ contains no points of $\partial P$ in its interior, all points in $\mathcal{E}[u,v]\cap P_{g}$ see all of $[u,v]$. By general position, $\mathcal{E}[u,v]$ either is left of $\overline{i\,j}$ and intersects $\overline{i\,j}$ in at most one point, or contains at least one point strictly right of $\overline{i\,j}$. We define $g^{*}$ as $i$ (or $j$) if $\mathcal{E}[u,v]$ contains $i$ (or $j$) and is left of $\overline{i\,j}$, or as the point in $\mathcal{E}[u,v]$ that maximizes the angle $\sphericalangle(i,g^{*},j)$, otherwise.
@@
Figure 6: Case 2 of Lemma˜5. The point $g^{*}$ that has $\sphericalangle(i,g^{*},j)=\alpha$ is in the visibility core $\mathcal{E}[u,v]$ (green), but not in $P_{g}$ (gray).
@@
Case 1: $\sphericalangle(i,g^{*},j)>\pi$. The point $g^{*}$ must lie strictly right of the edge $\overline{i\,j}$ and, in particular, cannot serve as an angle-maximising dominator. Instead, we construct the segment-intersecting dominator ($\mathbf{A}$) and show that it dominates $(g,[u,v])$. Because $\sphericalangle(u,g,v)\leq\pi$, we also have $\sphericalangle(i,g,j)\leq\pi$, which, together with $\sphericalangle(i,g^{*},j)>\pi$, implies that $\overline{i\,j}$ lies in $P_{g}$. In particular, $g^{*}$ also lies in $P_{g}$, with $\overline{i\,j}$ separating $g$ and $g^{*}$. By definition, both $g$ and $g^{*}$ lie in $\mathcal{E}[u,v]$. Since $\mathcal{E}[u,v]$ is convex, and $\overline{i\,j}$ separates $g$ and $g^{*}$, $\mathcal{E}[u,v]$ must intersect $\overline{i\,j}$. Moreover, every edge of $\mathcal{E}[u,v]$ supports an edge in [i1,j+1][i{-}1,j{+}1], and as no three vertices of $P$ are collinear, there must exist an intersection point in the interior of $\overline{i\,j}$.
Any point of intersection between $\overline{i\,j}$ and $\mathcal{E}[u,v]$ lies in both $\mathcal{E}[u,v]$ and $P_{g}$, and thus sees all of $[u,v]$. It follows that the last point of intersection $g^{\prime}$ between $\overline{i\,j}$ and $\mathcal{E}[u,v]$ sees $[u,v]$ and so the segment-intersecting dominator $(g^{\prime},[u_{\max},v_{\max}])$ corresponding to $(i,j)$ is well-defined and it dominates $(g,[u,v])$.
@@
Case 2: $g^{*}=i$ or $g^{*}=j$ or $\sphericalangle(i,g^{*},j)\leq\pi$. If we can show that $g^{*}$ is in $P_{g}$, then it lies in $P_{g}\cap\mathcal{E}[u,v]$ and thus sees all of $[u,v]$. So, the angle-maximizing dominator $(g^{*},[u_{\max},v_{\max}])$ ($\mathbf{B}$) corresponding to $(i,j)$ dominates $(g,[u,v])$. If $g^{*}=i$ or $g^{*}=j$ then $g^{*}$ in $P_{g}$, concluding the proof. So, assume that $g^{*}\neq i,j$ and for the sake of contradiction that $g^{*}\not\in P_{g}$, refer to Figure˜6. Let $\alpha=\sphericalangle(i,g^{*},j)$. The points $x\in\mathbb{R}^{2}$ such that $\sphericalangle(i,x,j)=\alpha$ form a circular arc passing through $i$ and $j$. As no other point in $\mathcal{E}[u,v]$ has larger angle than $\alpha$, and $\mathcal{E}[u,v]$ is convex, the tangent $t$ of this circular arc at $g^{*}$ separates the plane into two half-planes: a closed half-plane $H_{g}$ that contains $\mathcal{E}[u,v]$ and an open half-plane $H_{i,j}$ containing the points $i$ and $j$ (if $i$ or $j$ lies on $t$ then they coincide with $g^{*}$, contradicting our assumptions).

### v4

@@ 3.1 Upper bound
We begin with a concept that appears in several previous approaches [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover]. For any $u\in\partial P$, let $v\in[u,2n+1)$ be the maximal value such that there exists a guard $(g,[u,v])$. This defines a function $\textnormal{{next}}:\partial P\to\partial P$ mapping $u$ to $v$. Let $k$ denote the minimum number of guards required to cover $\partial P$. If we choose any point $x\in\partial P$ and apply next a total of $k+1$ times, the resulting point lies beyond $x+n$. Our goal is to compute a set $X\subseteq\partial P$ such that for at least one $x\in X$ there exists an optimal solution of $k$ guards containing a guard $(g,[x,v])$. If we recursively apply the next function $k$ times to all points in $X$, this yields an optimal solution using $k$ guards. Hence, if $X$ has linear size and if we can evaluate next in $O(\log n)$ time, we obtain an $O(kn\log n)$-time algorithm.
Classifying guards.Section˜4 introduces the structural results to compute $X$ and the next function. We define three collections of highly structured guards such that every guard is dominated by one of these three types. Specifically, we partition the set of all guards into good and bad guards. Intuitively, a guard $(g,[u,v])$ is good if the angle $\sphericalangle(u,g,v)$ is at most $\pi$, and bad otherwise. We then define three dominators: the good, the bad, and the ugly.
@@
We can now provide the high-level intuition for why $\textnormal{{next}}(x)$ can be evaluated in logarithmic time. If we precompute all good and bad dominators $(g,[u,v])$, then we can store their corresponding intervals $[u,v]$ in a segment tree $T$. For any $x\in\partial P$, we perform a stabbing query on $T$ and we return the maximal right endpoint $r$ among the intervals stabbed by $x$. Let $v=\textnormal{{next}}(x)$. If there exists a guard $(g,[u,v])$ dominated by a good or bad dominator, then $v=r$. We then compute an ugly dominator for $x$ on the fly and compare its result to $r$ to obtain the correct value of $\textnormal{{next}}(x)$.
@@
Defining a set $X$.  The reduced good dominators $\mathcal{D}$ are not only essential for computing next, but they also induce the set $X$. Specifically, we prove that there always exists an optimal solution Opt to the Contiguous Art Gallery problem such that at least one guard $(g,[u,v])\in\textsc{Opt}$ either belongs to $\mathcal{D}$ or has $u$ as a vertex of $P$. We define $X$ as the set of all vertices of $P$, together with all points $u\in\partial P$ such that there is a reduced good dominator $(g,[u,v])$.
Sliding windows.  We use a sliding window over indices $(i,j)$ to compute (a superset of) $\mathcal{D}$. The remainder of our algorithms rely on a different but very specific sliding window:
Definition 3.1.
@@
Computing $\sigma$ and the reduced dominators.  In Section˜6 we compute linear-size sets $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ where $\mathcal{D}\subseteq\mathcal{D}^{\prime}$ and $\mathcal{B}\subseteq\mathcal{B}_{\sigma}$. We first compute $\mathcal{D}^{\prime}$ using a sliding-window algorithm. Note that we cannot use a conforming sliding sequence $\sigma$ for this task, as a reduced good dominator may be strictly dominated by a guard that is not a good dominator. Consequently, the defining pair $(i,j)$ for a guard $(g,[u,v])\in\mathcal{D}$ need not belong to $\sigma$.
@@
For all subsequent algorithms, we use a conforming sliding sequence $\sigma$, which we compute in an online fashion starting from $(i,j)=(1,1)$ and repeatedly incrementing either $i$ or $j$. The choice of which index to increment follows a simple rule: while there exists a guard $(g,[u,v])$ such that $i,j+1]\subseteq[u,v]$, we increment $j$ and add the new pair $(i,j)$ to $\sigma$; otherwise, we increment $i$ and add the new pair $(i,j)$ to $\sigma$. By the result of [Chan2019Time-Windowed], we can maintain $\mathcal{E}[i-1,j+2]$ as a balanced binary tree while incrementing $i$ and $j$. We show that, given $i$, $j$, and $\mathcal{E}[i-1,j+2]$, this decision can be made in $O(\log n)$ time, yielding an overall $O(n\log n)$ algorithm for constructing $\sigma$. From $\sigma$, we derive a linear-size set $\mathcal{B}_{\sigma}$ containing $\mathcal{B}$. We store the visibility of guards in $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ in a segment tree $T$. Thus, we may obtain for any point $x\in\partial P$ the guard $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$ that has $x\in[u,v]$ and maximizes $v$.
@@
An $O(kn\log n)$ algorithm.  In Section˜7, we consider the set $X$ and iterate over all $(i,j)\in\sigma$. For each $x\in X$, we perform a stabbing query on $T$ to find the maximum value $r$ such that there exists a guard $(g,[x,r])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$. The value $\textnormal{{next}}(x)$ differs from $r$ only if it is realised by an ugly dominator. The value $\lfloor r\rfloor$ serves as a lower bound for the index $j$ such that $\textnormal{{next}}(x)\in(j,j+1]$, and by definition $(i,j)\in\sigma$. These observations yield an algorithm that iterates over $\sigma$ and computes $\textnormal{{next}}(x)$ for all $x\in X$ in $O(n\log n)$ total time. Repeating this process $k$ times produces an optimal solution in $O(kn\log n)$ time. This is already a major improvement over the state-of-the-art [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover] and serves as a crucial subroutine for our $O(n\log n)$-time algorithm.
@@
An $O(n\log n)$ algorithm.  We represent next as a piecewise, possibly degenerate, Möbius transformation. This approach was first proposed in [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover], where the resulting function consisted of $O(n^{3})$ pieces and was computed in $\tilde{O}(n^{5})$ time. In contrast, we use our dominators to define a function with only $O(n)$ pieces, which can be computed in $O(n\log n)$ time. Specifically, in Section˜8, we partition the boundary $\partial P$ into intervals $I$ such that either:
@@
3.2 Lower bound.
In Section˜10, we reduce from Set Disjointness, where we are given two sets $A,B\subseteq[0,n^{3}]$, each of size $n$, and must decide whether $A\cap B=\emptyset$. While in the wordRAM-model there is a deterministic $O(n\log\log n)$-time algorithm for solving this problem (via sorting [Han2002Deterministic]), in comparison-based models of computation such as the realRAM-model, it has a worst-case lower bound of $\Omega(n\log n)$, even if $A$ is already sorted [YaoLowerBound]. Given $A,B\subseteq[1,n^{3}]$ with $A$ sorted, we can in linear time construct a simple polygon $P$ such that $A\cap B=\emptyset$ if and only if $P$ can be guarded with $2|B|+2$ guards (see Figure˜2).
We begin with a triangle $T$ whose top edges, $t_{1}:$ $(0,0)$ to $(\frac{n^{3}}{2},\frac{n^{3}}{2})$ and $t_{2}:$ $(\frac{n^{3}}{2},\frac{n^{3}}{2})$ to $(n^{3},0)$, represent the interval $0,n^{3}]$. We construct the polygon in cyclical clockwise order, starting from $(0,0)$. For each of the increasing values $a_{i}\in A$, we make an indent in $t_{1}\cup t_{2}$ excluding the point corresponding to $a_{i}$ from $P$. Next, we partition a small segment on the bottom edge ($(n^{3},0)$ to $(0,0)$) into $|B|$ intervals $I_{j}$. For each $b_{j}\in B$, we introduce a pocket: two consecutive indentations of $I_{j}$ such that a guard can see $I_{j}$ if and only if it stands at the point of $t_{1}$ or $t_{2}$ corresponding to $b_{j}$. Hence, there exists a guard that can cover $I_{j}$ if and only if $b_{j}\notin A$. Finally, we place a blocker between consecutive intervals $I_{j}$ and $I_{j+1}$ such that each $I_{j}$ and each blocker requires a unique guard. The resulting polygon $P$ (specified by its vertices along the polygon boundary) can be constructed in linear time and guarded with $2|B|+2$ contiguous guards if and only if $A\cap B=\emptyset$, establishing an $\Omega(n\log n)$ lower bound for the Contiguous Art Gallery problem.
@@ 4 A combinatorial classification of guards
We first partition the set of all guards into good and bad guards. In Figure˜1, $g_{2}$, $g_{3}$, and $g_{4}$ are good, while $g_{1}$ and $g_{5}$ are bad.
Definition 4.1.
@@
We will define three collections of guards called dominators: the good, the bad, and the ugly. We prove that all good guards are dominated by a good dominator, which will induce a set $X$ of $O(n)$ points along $\partial P$ such that there exists an optimal solution that includes a guard $(g,[x,v])$ for $x\in X$. We prove that all bad guards are dominated by either a bad or an ugly dominator (see Figure 3). To this end, we first make some observations:
Lemma 4.2.
@@
Proof 4.3.
By maximality of $v$, $g\neq v$ and there exists an $\varepsilon^{*}>0$ such that $\forall\varepsilon\in(0,\varepsilon^{*}]$ the shortest path from $g$ to $v+\varepsilon$ visits a reflex vertex $x$. If $x=v$ then $v$ is a reflex vertex of $P$. Otherwise, $\overline{v\,g}$ contains $x$ in its interior.
Lemma 4.4.
@@
Proof 4.5.
For a guard that is inclusion-wise maximal it cannot be that $g=u$ or $g=v$. That the shortest path from $u$ to $v$ is left-turning follows from the fact that $g$ can see $u$ and $v$ and $\sphericalangle(u,g,v)>\pi$. If $u$ (respectively $v$) is not a vertex of $P$, then Lemma˜4.2 implies that the edge $\overline{u\,g}$ (respectively $\overline{v\,g}$) contains a reflex vertex of $P$ in its interior, which must appear as an interior vertex on the shortest path from $u$ to $v$.
If only $v$ is maximal, then it again cannot be that $g=v$, but it may be that $g=u$. In this case the shortest path from $u$ to $v$ is the segment $\overline{u\,v}$, which is also left-turning. If $v$ is not a vertex of $P$, Lemma˜4.2 still implies there is at least one vertex on this shortest path.
Figure 3: Flowchart illustrating how a guard is dominated. We prove the green block in Lemma˜4.9 and the red block in Lemma˜4.14.
The good dominators.  We show that there exists a set $\mathcal{D}$ of $O(n)$ guards, called the reduced good dominators, such that every good guard is dominated by one of them. We begin by defining two categories of good dominators, each of size $O(n^{2})$.
{observation}
@@
Proof 4.6.
For any fixed point $g\in P$, the set of points in $P$ visible from $g$ forms the visibility polygon of $g$. Its intersection with $\partial P$ induces a fixed set of pairwise disjoint maximal chains, and only one of these chains can contain $[i,j]$.
Definition 4.7.
@@ 1. $\mathbf{A}$
    (segment-intersecting): See Figure˜5. If $\mathcal{E}[i-1,j+1]$ intersects the interior of $\overline{i\,j}$, we define the guard $(g,[u_{\text{max}},v_{\text{max}}])$, where $g$ is the last point of $\overline{i\,j}$ intersecting $\mathcal{E}[i-1,j+1]$, and $u_{\text{max}},v_{\text{max}}]\subseteq[i-1,j+1]$ is the maximal chain containing $[i,j]$ that is visible to $g$.
@@ 2. $\mathbf{B}$
    (angle-maximizing): See Figure˜5. If $\mathcal{E}[i-1,j+1]$ lies left of $\overline{i\,j}$, we define the guard $(g,[u_{\text{max}},v_{\text{max}}])$. We define $g$ as $i$ (or $j$) if $\mathcal{E}[i-1,j+1]$ intersects $i$ (or $j$). Otherwise, $g$ is the point in $\mathcal{E}[i-1,j+1]$ that maximizes the angle $\sphericalangle(i,g,j)$. We define $u_{\text{max}},v_{\text{max}}]\subseteq[i-1,j+1]$ as its maximal visible chain containing $[i,j]$.
@@
The guard $(g,[u,v])$ is good and thus $g\neq u,v$ and $\sphericalangle(v,g,u)\leq\pi$ by Definition˜4.7. We prove that a good dominator (and thus a reduced good dominator) dominates $(g,[u,v])$. Let $u\in[i-1,i)$ and $v\in(j,j+1]$. If $i=j$ or i1=ji-1=j, then the dominator for $(i,j)$, which is the guard (i,[i1,j+1])(i,[i-1,j+1]), dominates $(g,[u,v])$. So, assume $i<j$.
@@
We define by $P_{g}$ the polygon formed by the edges $\overline{v\,g}$, $\overline{g\,u}$ and $[u,v]$. Since $(\overline{u\,g},\overline{v\,g})$ forms a convex wedge, and $P_{g}$ contains no points of $\partial P$ in its interior, all points in $\mathcal{E}[u,v]\cap P_{g}$ see all of $[u,v]$. By general position, $\mathcal{E}[u,v]$ either is left of $\overline{i\,j}$ and intersects $\overline{i\,j}$ in at most one point, or contains at least one point strictly right of $\overline{i\,j}$. We define $g^{*}$ as $i$ (or $j$) if $\mathcal{E}[u,v]$ contains $i$ (or $j$) and is left of $\overline{i\,j}$, or as the point in $\mathcal{E}[u,v]$ maximizing $\sphericalangle(i,g^{*},j)$, otherwise.
@@
Figure 6: Case 2 of Lemma˜4.9. The point $g^{*}$ that has $\sphericalangle(i,g^{*},j)=\alpha$ is in the visibility core $\mathcal{E}[u,v]$ (green), but not in $P_{g}$ (gray).
@@
Case 1: $\sphericalangle(i,g^{*},j)>\pi$. The point $g^{*}$ must lie strictly right of the edge $\overline{i\,j}$ and, in particular, cannot serve as an angle-maximising dominator. Instead, we construct the segment-intersecting dominator ($\mathbf{A}$) and show that it dominates $(g,[u,v])$. Because $\sphericalangle(u,g,v)\leq\pi$, we also have $\sphericalangle(i,g,j)\leq\pi$, which, together with $\sphericalangle(i,g^{*},j)>\pi$, implies that $\overline{i\,j}$ lies in $P_{g}$. In particular, $g^{*}$ also lies in $P_{g}$, with $\overline{i\,j}$ separating $g$ and $g^{*}$. By definition, both $g$ and $g^{*}$ lie in $\mathcal{E}[u,v]$. Since $\mathcal{E}[u,v]$ is convex, and $\overline{i\,j}$ separates $g$ and $g^{*}$, $\mathcal{E}[u,v]$ must intersect $\overline{i\,j}$. Moreover, every edge of $\mathcal{E}[u,v]$ supports an edge in [i1,j+1][i{-}1,j{+}1], and as no three vertices of $P$ are collinear, there must exist an intersection point in the interior of $\overline{i\,j}$. Any intersection between $\overline{i\,j}$ and $\mathcal{E}[u,v]$ lies in both $\mathcal{E}[u,v]$ and $P_{g}$, and thus sees all of $[u,v]$. It follows that the last point of intersection $g^{\prime}$ between $\overline{i\,j}$ and $\mathcal{E}[u,v]$ sees $[u,v]$ and so the segment-intersecting dominator $(g^{\prime},[u_{\max},v_{\max}])$ corresponding to $(i,j)$ is well-defined and it dominates $(g,[u,v])$.
@@
Case 2: $g^{*}=i$ or $g^{*}=j$ or $\sphericalangle(i,g^{*},j)\leq\pi$. If we can show that $g^{*}$ is in $P_{g}$, then it lies in $P_{g}\cap\mathcal{E}[u,v]$ and thus sees all of $[u,v]$. So, the angle-maximizing dominator $(g^{*},[u_{\max},v_{\max}])$ ($\mathbf{B}$) corresponding to $(i,j)$ dominates $(g,[u,v])$. If $g^{*}=i$ or $g^{*}=j$ then $g^{*}$ in $P_{g}$, concluding the proof. So, assume that $g^{*}\neq i,j$ and for the sake of contradiction that $g^{*}\not\in P_{g}$, refer to Figure˜6. Let $\alpha=\sphericalangle(i,g^{*},j)$. The points $x\in\mathbb{R}^{2}$ such that $\sphericalangle(i,x,j)=\alpha$ form a circular arc passing through $i$ and $j$. As no other point in $\mathcal{E}[u,v]$ has larger angle than $\alpha$, and $\mathcal{E}[u,v]$ is convex, the tangent $t$ of this circular arc at $g^{*}$ separates the plane into two half-planes: a closed half-plane $H_{g}$ that contains $\mathcal{E}[u,v]$ and an open half-plane $H_{i,j}$ containing the points $i$ and $j$ (if $i$ or $j$ lies on $t$ then they coincide with $g^{*}$, a contradiction).